produces a product with fixed costs of $54,100 and variable cost of $2.90 per unit. The $24,000 profit and believes it can sell 11,000 units of the product. Round your answer to 2 decimal places ed
60 108 $144,000 a. What is the break-even point in dollars? In units? b. To obtain a profit of $48,000, what must the sales be in dollars? In units? c. If the sales price increases to $72 and variable costs do not change, what is the new break-even point in dollars? In units? K Prey6 of 8 Next>
Answer #1
answer 1
target price =
Fixed cost 54100
Variable cost 11000*2.9 31900
Desired profit = 24000
total sales price 110000
Unit sold = 11000
Unit sales price =110000/11000 10 per unit
answer 2
Sales price 60
Contribution margin 10%
Fixed cost 144000
Therefore contribution margin =60*10% 6
variable cost =60-6 = 54 54
a Break even point in unit =Fixed cost / Contribution margin per unit
=144000/6 24000 unit
Break even point in Dollar =Fixed cost / Contribution margin %
=144000/10% 1440000
b Sales in Dollar =(Fixed cost+ 48000)/Contribution margin %
=(144000+ 48000)/10% 1920000
Sales in unit = 1920000/60 32000
c Increase in sales price = 72
revised contribution margin =72-54 18
revised contribution margin %=18/72 25%
Break even point in unit =Fixed cost / Contribution margin per unit
=144000/18 8000 unit
Break even point in Dollar =Fixed cost / Contribution margin %
=144000/25% 576000